OpenStudy (anonymous):

Find the slope of the line that passes through the pair of points (–1, –9) and (–3, 10)

OpenStudy (blurbendy):

10 + 9 = 19 -3 + 1 = -2 slope = 19/-2

OpenStudy (anonymous):

thanks i knew it was simple i just haven't done thins in a while.

OpenStudy (blurbendy):

np!